# Towards Verifiable Text Generation with Evolving Memory and Self-Reflection

### 14 Dec 2023 v1

@@ 3.1 Baselines
For an equitable comparison, we have selected four best-performing baseline methodologies as proposed in ALCE Gao et al. (2023b), which include Vanilla, Summ, Snippet and Rerank. Each of these methods incorporates multiple demonstrations within the initial prompt to facilitate the process of generating responses.
@@
This methodology entails a two-stage process. Initially, the LLM generates four distinct responses using high temperature. Thereafter, each response undergoes an evaluation for citation recall. The response with the highest citation recall score is then chosen as the final output.
3.2 Datasets and Evaluation
We assess the effectiveness of our methods on five datasets across three knowledge-intensive tasks, as detailed in Table 4 of Appendix A. For all datasets, our evaluation criteria encompass both the answer correctness and citation quality of model outputs.
Multihop QA
Long-form QA
aims to generate detailed answers to complex questions Fan et al. (2019a, b), we utilize the ASQA Stelmakh et al. (2022) and ELI5 Fan et al. (2019b) datasets. ASQA focuses on ambiguous questions requiring comprehensive answers covering multiple interpretations. ELI5, on the other hand, deals with complex questions demanding lengthy, in-depth answers backed by multiple documents.
Open-domain QA
requires leveraging external knowledge for answering questions. Our datasets of choice are NQ Kwiatkowski et al. (2019) and WebQ Berant et al. (2013).
Following the methodology of Sun et al. (2023), LLMs are prompted to generate the final answer, which is then compared with the reference answer using answer-level exact match (EM), and token-level F1.
Verifiability Evaluation
3.3 Implementation Details
To ascertain the generalizability of our method, we conducted experiments using different sizes of Large Language Models (LLMs). Specifically, we utilized Vicuna-13B-v1.5-16k Zheng et al. (2023) [^1] https://huggingface.co/lmsys/vicuna-13b-v1.5-16k and Text-Davinci-003 Ouyang et al. (2022) [^2] https://api.openai.com/v1/completions as of October 2023 for evaluation, respectively.
In alignment with prior research Trivedi et al. (2022); Jiang et al. (2023), we randomly subsampled 500 examples for Vicuna-13B-v1.5-16k and 200 examples for Text-Davinci-003 from each dataset due to the cost of running experiments. For the evaluation of verifiability and inference tasks in both the Rerank and Vtg models, we employed the TRUE [^3] https://huggingface.co/google/t5_xxl_true_nli_mixture model (Raffel et al., 2020), a T5-11B model fine-tuned on a collection of NLI datasets, to automatically examine whether the cited documents entail the model generation.
3.4 Main Results
In this section, we delineate our primary findings derived from extensive analyses conducted on five diverse datasets, as outlined in Tables 1 and 2. These investigations yield several key observations:
First, our proposed Vtg method consistently outperforms other approaches across various datasets and metrics when applied to LLMs of different sizes. Notably, Vtg enhances citation quality significantly, achieving a remarkable 20% and 9% improvement over our strongest competitor Rerank, using Text-Davinci-003 and Vicuna-13B-v1.5-16k, respectively. Furthermore, Vtg’s capacity for verifiable generation substantially enhances correctness, evidenced by an approximate 6% overall improvement compared to the leading baseline across both LLMs.
Second, among the evaluated baselines, Rerank stands out in citation quality, primarily due to its multiple sampling strategy that increases the likelihood of high-quality outputs. However, its performance in terms of correctness is inconsistent across different datasets. For instance, Rerank’s effectiveness is notably lower in ASQA’s EM using Text-Davinci-003 and ELI5’s Claim using vicuna-13B-v1.5-16k. In contrast, Vtg demonstrates consistent improvement across all metrics and datasets, highlighting its robustness and reliability.
Third, the comparison between two different LLMs, i.e., vicuna-13B-v1.5-16k and Text-Davinci-003, reveals intriguing distinctions. Methods like Summ and Snippet, which integrate a broader range of documents, achieve greater correctness with vicuna-13B-v1.5-16k but falter with Text-Davinci-003. This disparity can be attributed to Text-Davinci-003’s comparatively extensive knowledge base, as evidenced by the significant improvement in correctness when applying identical methodologies to both LLMs. Consequently, Summ and Snippet may introduce extraneous information in the context of Text-Davinci-003. Moreover, these methods underperform in citation quality with both LLMs, potentially due to the fact that their reliance on simplified document versions challenges the verification of document support for the claims.
Datasets Correct Citation
EM F1 Rec Pre F1
Vtg 25.60 23.27 55.36 49.59 52.32
-w/o QG 21.20 18.98 48.97 41.54 44.95
-w/o Verifier 21.00 17.72 33.09 20.07 24.98
-w/o Simplifier 24.40 22.23 46.38 29.47 36.04
Table 3: Ablation Study on Wikihop.
3.5 Ablation Study
In this section, we evaluate the impact of each component in our model using the Wikihop dataset. By removing components one by one, we observe their individual contributions to performance. “w/o QG” replaces query generation with direct retrieval using the original sentence. “w/o Verifier” excludes the two-tier verifier during generation, and “w/o Simplifier” removes the citation simplification.
Results in Table 3 show that removing any component decreases performance, highlighting their importance. Notably, the removal of the verifier leads to the most significant drop, as it results in the generation of sentences without external validation, potentially introducing hallucinations into the generation. On the other hand, omitting the simplifier has a less pronounced effect on correctness, mainly affecting citation quality through reduced precision. Removing QG limits the model’s ability to retrieve a broader range of relevant documents, leading to performance degradation.
3.6 Hyper-parameter Analysis
In this section, we examine the impact of various hyper-parameters on Vtg’s performance on ASQA dataset, with a focus on the number of documents retrieved per query $N$, the number of generated queries $M$, and the maximum trials $T$, as illustrated in Figure 3. We adjust these parameters one at a time to observe their individual effects. Our findings are as follows:
For the number of documents retrieved per query $N$, initially increasing $N$ enhances citation quality but reduces answer correctness. This trend suggests that a larger pool of documents provides more citation options for the LLM, but also introduces potentially noisy information, impacting model generation. When $N$ becomes very large, there’s a consistent decrease in both citation quality and answer correctness, likely due to the overwhelming amount of information confusing the model.
For the number of generated queries $M$, the optimal performance for both answer correctness and citation quality occurs at $M$=2. Beyond this, performance declines. When $M$ is small, increasing it broadens the knowledge scope for document retrieval, thereby improving outcomes. However, a larger $M$ can lead to less relevant queries, introducing noise and thus degrading model performance.
For the maximum trials $T$, a higher $T$ correlates with improved performance, as it allows the model more attempts to generate a sentence that passes verification. This approach provides the model with more chances for a successful generation. However, an increase in $T$ also results in higher token consumption, indicating a need to balance effectiveness with computational cost. Adjusting $T$ can help find an equilibrium between model performance and computational cost.
Figure 4: Token Consumption Comparison on NQ.
3.7 Token Consumption Analysis
This section presents a comparative analysis of token consumption by different methods on the NQ dataset, as detailed in Figure 4.
As we can find, Summ, Snippet, and Vanilla incur nearly identical token costs per query. Rerank, which samples responses four times, incurs the highest token cost, albeit yielding superior citation quality compared to other baselines. Contrarily, our method Vtg demonstrates a lower token cost than Rerank while significantly enhancing citation quality, thereby underscoring the efficiency of our approach. Furthermore, as discussed in Section 3.6, the token cost in our method can be modulated through the max trials hyperparameter, providing a versatile option for balancing performance with resource expenditure.
4 Related Work
4.1 Retrieval-augmented LLMs.
Retrieval-augmented LLMs aim to provide extra documents to the LLM, which has been proven useful in many knowledge-intensive tasks. Among the existing studies, Shi et al. (2023); Wang et al. (2023b); Zhang et al. (2023); Yu et al. (2023a, c) propose to retrieve only once at the beginning. Other works Qian et al. (2023); Yu et al. (2023b) propose to retrieve multiple times during generation, which offers flexibility of when and what to search. For example, Jiang et al. (2023) propose to retrieve when the generated sentence contains low confidkkence tokens. Ram et al. (2023) propose to refresh the retrieved document every n token, which is demonstrated to be more effective than retrieving only once. Wang et al. (2023a); Asai et al. (2023); Zhao et al. (2023) propose to retrieve only when the LLMs need to.
Among the existing studies, retrieve on-the-fly methods are the closest to ours. However, these methods do not provide referenced documents for the generated sentences, potentially reducing the reliability of the generated content.
4.2 Verifiable Text Generation
Verifiable Text Generation aims to generate content with supporting documents, which has been attracting attention in recent years. For example, Liu et al. (2023b); Qin et al. (2023); Nakano et al. focus on training LLMs to browse web pages and answer questions with evidence. Gao et al. (2023a) introduced the research-and-revision (RARR) method for retrieving evidence for LLM outputs. Li et al. (2023) incorporated knowledge graphs as an evidence source. Other works mainly focus on evaluation Liu et al. (2023a). For example, Rashkin et al. (2023) propose Attributable to Identified Sources (AIS) for human evaluation. Gao et al. (2022) define auto-AIS to approximate human AIS judgments. Gao et al. (2023b) propose ALCE to automate the evaluation of the citation quality. Min et al. (2023) proposed FactScore for evaluating the verifiability of generated facts.
However, these methods do not effectively address the focus-shifting problem. In this paper, we propose incorporating a long short-term memory to retain important documents in the long-term memory and current documents in the short-term memory.
5 Conclusion
In this paper, we introduce Vtg, a novel framework tailored to address the complexities of verifiable text generation. Central to its design is an evolving long short-term memory, which adaptively keeps both valuable documents and up-to-date documents. Through the integration of active retrieval mechanisms and diverse query generation, Vtg skillfully enhances both the precision and scope of the document retrieval process. The two-tier verifier coupled with an evidence finder facilitates a deeper analysis and reflection on the relationship between claims and citations. Extensive experiments on five datasets across three knowledge-intensive tasks verify the effectiveness of our method.
Appendix A Datasets and Settings
@@
Settings 2WikiMultihopQA ASQA ELI5 NQ WebQ
Ho et al. (2020) Stelmakh et al. (2022) Fan et al. (2019b) Kwiatkowski et al. (2019) Berant et al. (2013)
Dataset statistics
Task Multihop QA Long-form QA Long-form QA Open-domain QA Open-domain QA
#Examples (Vicuna) 500 500 500 500 500
#Examples (Davinci) 200 200 200 200 200
Evaluation settings
Correctness Metrics EM, Token-level F1, Pre, and Rec EM, Disambig-F1, ROUGE, DR Claim EM, F1 EM, F1
Citation Metrics Prec, Rec, F1
Retrieval settings
Corpus Wikipedia Wikipedia Wikipedia Wikipedia Wikipedia
Retriever DPR DPR DPR DPR DPR
Table 4: Statistics and experimental settings of different tasks/datasets.
Appendix B Prompts
@@
Given the claim: claim, please generate up to num questions that can help verify the claim with the following constraints: 1. You should output no more than three questions. 2. The generated questions should be diverse and focus on different aspects of the given question. Generated questions:
Prompt B.4: Summ Instruction
Summarize the following document within 50 words with the question of interest {Question} Return "irrelevant" if the document is “irrelevant" to the question. Try to keep all the important dates, numbers, and names. Title: {Title} Text: {Text} Summary:
Prompt B.5: Snippet Instruction
Given the following passage and the question {Question}, extract a useful span from the passage that can answer the question. Resolve all the coreference issues to make the extracted span understandable and standalone. If the passage is not helpful for answering the question, return “irrelevant". If there are multiple spans, merge them and only output one paragraph. Title: {Title} Text: {Text} Extracted span:

### 11 Jan 2024 v2

@@ Appendix A Baselines
For an equitable comparison, we have selected four best-performing baseline methodologies as proposed in ALCE Gao et al. (2023b), which include Vanilla, Summ, Snippet and Rerank. Each of these methods incorporates multiple demonstrations within the initial prompt to facilitate the process of generating responses. Following ALCE Gao et al. (2023b), we set $k=5$ and $K=10$ in our experiment.
@@
This methodology entails a two-stage process. Initially, the LLM generates four distinct responses based on the top-$k$ ranked documents using high temperature. Thereafter, each response undergoes an evaluation for citation recall. The response with the highest citation recall score is then chosen as the final output.
Appendix B Datasets and Settings
@@
Citation Recall.
Citation recall for each claim in the model’s response is computed individually as either 0 or 1 and then averaged across all claims in the response. A claim’s citation recall is 1 if at least one citation exists and the concatenated citations entail the claim according to an NLI model, which outputs 1 for entailment.
Citation Precision.
Citation precision for each citation in the model’s response is computed individually as either 0 or 1 and then averaged across all citations in the response. The precision score for a citation is 1 if the associated claim has a citation recall of 1 and the citation is not irrelevant; otherwise, it’s 0. A citation is deemed irrelevant if: (a) the citation alone cannot substantiate the claim, and (b) omitting the citation doesn’t impact the remaining citations’ ability to support the claim.
Appendix C Hyper-parameters
The detailed hyper-parameters used in Vtg for Text-Davinci-003 and Vicuna-13B-v1.5-16k are shown in Table 6 and Table 7, respectively.
Appendix D Algorithm
The algorithm procedural of Vtg is shown in Algorithm 1
Appendix E Prompts
The prompts used in our experiments are listed as follows. It’s worth noting that the prompts for Vanilla and Rerank are identical, so we only present the one for Vanilla.
@@ Appendix E Prompts
Given the original question: {Question}. The context is as follows: {Context}. The claim is: {Claim}. Please generate up to {qg_num} questions that can help verify the claim with the following constraints: 1. You should output no more than {qg_num} questions. 2. The generated questions should be diverse and focus on different aspects of the given claim. Generated questions:
Prompt for Vanilla
Instruction: Write a high-quality answer for the given question using only the provided search results and cite them properly using [1][2][3]. Question: {Question} Document: {Document} Answer:
Prompt for Summ
## Step 1: First Summarize the documents Summarize the following document within 50 words with the question of interest {Question} Return "irrelevant" if the document is “irrelevant" to the question. Try to keep all the important dates, numbers, and names. Title: {Title} Text: {Text} Summary: ## Step 2: Generate the response based on the summary Instruction: Write a high-quality answer for the given question using only the provided search results and cite them properly using [1][2][3]. Question: {Question} Document: {Document} Answer:
Prompt for Snippet
## Step 1: First extract relevant snippet from the documents Given the following passage and the question {Question}, extract a useful span from the passage that can answer the question. Resolve all the coreference issues to make the extracted span understandable and standalone. If the passage is not helpful for answering the question, return “irrelevant". If there are multiple spans, merge them and only output one paragraph. Title: {Title} Text: {Text} Extracted span: ## Step 2: Generate the response based on the snippet Instruction: Write a high-quality answer for the given question using only the provided search results and cite them properly using [1][2][3]. Question: {Question} Document: {Document} Answer: